OpenStudy (anonymous):

Find (f • g)(x) for f(x) = x + 7 and g(x) = x – 3

OpenStudy (australopithecus):

so just sub g into f so you want f(g(x))

OpenStudy (australopithecus):

that dot means to sub the one function into the other function

OpenStudy (australopithecus):

do you just want me to show you or do you follow?

hero (hero):

I'm pretty sure that dot means multiplication: (f • g)(x) = f(x) • g(x) = (x+7) • (x-3) Don't confuse (f • g)(x) with (f o g)(x)

OpenStudy (australopithecus):

oh yeah you are right hero

OpenStudy (anonymous):

still kinda confused

hero (hero):

Why? It's pretty obvious to multiply (x+7)(x-3)

OpenStudy (australopithecus):

yeah unless it the dot is hollow then it is composition of functions

OpenStudy (australopithecus):

I just dont like the use of dots for multiplication because of their use in vectors

OpenStudy (australopithecus):

or linear algebra rather

OpenStudy (anonymous):

do foil?

OpenStudy (anonymous):

or just 7x3?

hero (hero):

Do distributive rule (x+7)(x-3) = x(x-3) + 7(x-3) = x^2 - 3x + 7x - 21 = x^2 +4x - 21

hero (hero):

FOIL is an acronym not a mathematical property

OpenStudy (anonymous):

ok